# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

### v2

@@ I.3 Algorithm Unrolling with an Undirected Temporal Graph
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{n}^{\tau}$ is also similar to eq. 18 as follows:
${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\bm{\gamma}}_{n}^{\tau+1}={\bm{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix F.2. The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving eq. 78 follows the setup in Appendix F.1.
@@
Dataset PEMS03 METR-LA
$2\sigma_{\text{RMSE}}$ 0.05 0.04
$2\sigma_{\text{MAE}}$ 0.02 0.02
$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04

### v3

@@ I.3 Algorithm Unrolling with an Undirected Temporal Graph
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (78)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (79)
$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (80)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{n}^{\tau}$ is also similar to Equation 18 as follows:
${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\boldsymbol{\gamma}}_{n}^{\tau+1}={\boldsymbol{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (81)
The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix F.2. The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving Equation 80 follows the setup in Appendix F.1.
@@
Dataset PEMS03 METR-LA PEMS08
$2\sigma_{\text{RMSE}}$ 0.05 0.04 0.04
$2\sigma_{\text{MAE}}$ 0.02 0.02 0.02
$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04 0.04